Grade 12th passPhysical Chemistry

Calculate minimum uncertainty in the speed of ball mass 500g in 5 micro meter of a certain point on the ball

Chitra
7 Years agoGrade 12th pass

1 Answer

Arun
7 Years ago
Dear student
 
deltaV = h / 2m * deltax = 1.05 * 10^-34 / (2* 0.5 * 1 * 10^-6) = 1.05 * 10^-28 m/sec
 
hope it helps
 
Regards
Arun (askIITians forum expert)